The decomposition reaction :-


4HNO3(g)  <-> 4NO2(g) +2H2O(g)+O2(g) .


Is started with pure HNO3 if {P} is the total pressure at equilibrium,then prove that:-


Kp=1024(Po2*Po2*Po2*Po2*Po2*Po2*Po2)/{(p-7Po2)(p-7Po2)(p-7Po2)(p-7Po2)


EXPLAIN HOW???


 

Grade:11

1 Answers

Mehvash
5 years ago
4HNO3=4NO2+2H2O+O2
let partial pressure be p
Ptotal=P(HNO3)+P(H2O)+P(O2)
Initially for HNO3 pressure =p
and NO2,H2O,O2= 0
at equilibrium: for HNO3 pressue=P- (PHNO3+PH2O+PO2)
4HNO3= 4NO2+2H2O+O2
Initially:P    0        0        0
equilibrium:p-(phno3+ph2o+po2)=phno3  ph2o   po2
(in terms of po2), PNO2=4PO2
PH2O=2PO2
KP=P[NO2]4 P[H2O]2 P[O2] / P[HNO3]
KP=[4PO24][2PO2]2[O2] / [P-7PO2]4
KP=1024[PO2]7 / [P-7PO2]4